OpenStudy (anonymous):

how do I add 6g/2g^4+5-g/2g^4

OpenStudy (anonymous):

I got 5g+5/2g^4

OpenStudy (anonymous):

Solve for g over the real numbers: 5/2 (g^5+2) = 0 Multiply both sides by a constant to simplify the equation. Multiply both sides by 2/5: g^5+2 = 0 Isolate terms with g to the left hand side. Subtract 2 from both sides: g^5 = -2 Elminate the power on the left hand side. Take 5^th roots of both sides: Answer: | | g = -2^(1/5)

OpenStudy (anonymous):

That can't be the final answer